# SSHPool: The Separated Subgraph-based Hierarchical Pooling

### v1

@@ Abstract
In this paper, we develop a novel local graph pooling method, namely the Separated Subgraph-based Hierarchical Pooling (SSHPool), for graph classification. To this end, we commence by assigning the nodes of a sample graph into different clusters, resulting in a family of separated subgraphs. We individually employ a local graph convolution units as the local structure to further compress each subgraph into a coarsened node, transforming the original graph into a coarsened graph. Since these subgraphs are separated by different clusters and the structural information cannot be propagated between them, the local convolution operation can significantly avoid the over-smoothing problem arising in most existing Graph Neural Networks (GNNs). By hierarchically performing the proposed procedures on the resulting coarsened graph, the proposed SSHPool can effectively extract the hierarchical global feature of the original graph structure, encapsulating rich intrinsic structural characteristics. Furthermore, we develop an end-to-end GNN framework associated with the proposed SSHPool module for graph classification. Experimental results demonstrate the superior performance of the proposed model on real-world datasets, significantly outperforming state-of-the-art GNN methods in terms of the classification accuracies.
@@ I Introduction
In recent years, deep learning methods such as the Convolutional Neural Network (CNN) has experienced a rapid developing process. Many effective GNN based models have been applied to multiple research fields, including image segmentation [1], image classification [2], natural language processing [3], etc. The great success of the CNN encourages researchers to further explore employing it on graph structure data [4], that usually has a complex structure and is a typical instance of irregular data in non-Euclidean space. Unlike the regular data such as images and vectors, the graph data requires more effective models, that can not only analyze the attributes of the elements in the graph structure but also figure out the structural relationships between the elements. Thus, it is difficult to directly employ the traditional CNN for analyzing the graph data.
@@
To effectively deal with the graph data, some researchers have tried to enhance the ability of CNN models for irregular grid structures of graphs. Meanwhile, the Graph Neural Networks (GNNs) [5] are proposed to accomplish graph-based tasks. By integrating the CNN and the GNN methods together, the Graph Convolutional Network (GCN) [6] has been rapidly developed, and its related approaches [7, 8] have achieved outstanding performance in capturing the effective features of the graph data. For example, the Deep Graph Convolution Neural Network (DGCNN) [8] has leveraged the graph convolution layers with the SortPooling module to learn semantic characteristics. Based on the DGCNN models, some researchers further explore its application for different downstream tasks [9, 10, 11]. To accomplish the classification task, the Hierarchical Graph Convolutional Network (HGCN) [12] has adopted the hierarchical graph convolution to aggregate structurally similar nodes, exhibiting its strong empirical performance on some specific downstream tasks (e.g., node classification). Furthermore, by associating with the attention mechanism, some derivative GNN models, such as the Graph Transformer Network (GTN) [13] and the Spatio-Temporal Graph Transformer Network (STGTN) [14], have be developed and further improve the capability of extracting complicated graph features. These extended GNN-based models provide powerful tools for graph data analysis. However, when facing the graph representation learning task, there is usually a demand for these aforementioned models, i.e., they need to abstract features from original graphs with less redundant information [15]. To this end, the downsampled filtering method, namely the graph pooling operation, is highlighted.
@@
Similarly to the traditional pooling operation defined for the CNN models, the graph pooling operation proposed for the GNN models plays an important role in the sampling process. The main purpose of the graph pooling is to downsize the graph structure and make the representation learning efficient [16]. For existing works, the hierarchical scheme in the GNN models has significantly affected the definitions of novel graph pooling operations, and the researchers have focused more on the enhancement of the GNN models associated with the idea of hierarchical pooling [17, 18, 19, 20, 21]. Specifically, for the multi-layer framework of the GNN models, the original input graph can be gradually compressed into a family of coarsened graphs with shrinking sizes, so that we can get the global feature of the original graph through the coarsened graphs. According to the arrangement of nodes involved in the assignment, there are two specific types of pooling methods for graphs, namely the global pooling and the local pooling. The descriptions of both the methods are shown in Figure 1.
@@
Figure 1: Descriptions of two kinds of pooling methods. Both the global and the local pooling methods convert the original input graph into a global feature. On the left side, the global pooling tends to directly sum up or average the representations of all nodes over the whole graph structure. By contrast, on the right side, the local pooling tends to assign the nodes into different clusters, and then hierarchically aggregates the nodes from the same cluster into a coarsened node, resulting in a family of layer-wise coarsened graphs.
@@
Generally speaking, the global pooling is a common methodology in the pooling operation [22, 23]. Global operations, such as the summing pooling and the averaging pooling, can directly compute the global representation of a graph by simply summing up or averaging all the node representations over the whole graph [24]. However, since the global pooling neglects the feature distribution associated with different nodes, it cannot identify the structural differences between the nodes. Thus, the global pooling can not effectively learn the topological information residing on the local nodes and may be unsuitable for a hierarchical structure. To overcome the shortcoming, the local pooling that focuses more on the features of local nodes is developed. Based on this strategy, a number of local pooling operations, namely the hierarchical pooling operations, based on hierarchical structures are developed. For instance, Ying et al. [17] have proposed a differentiable graph pooling (DiffPool) module to generate the hierarchical representations from original graphs. The core idea is to learn the layer-wise node assignment matrices, that assign the nodes to different clusters and then hierarchically aggregate the graph of the current layer into a coarsened graph for the next layer, by compressing the nodes belong to the same cluster into a new coarsened node. However, since the the associated node assignment matrix is computed based on the soft assignment procedure, i.e., each node may be simultaneously assigned to multiple clusters associated with different probabilities. For the DiffPool, the receptive field of each local patch (i.e., the cluster) may influence that of other patches. Moreover, the coarsened node representation of the DiffPool is computed by directly summing up the node representations belong to the same cluster, the DiffPool fails to identify the differences between different nodes. To overcome the aforementioned problems, Ranjan et al. [18] have designed an adaptive structure aware pooling (ASAPool) method that can select the top scoring clusters after assigning the nodes of the the input graph during the pooling process. Liu et al. [21] have developed an attention-based differentiable pooling (ABDPool) operation, that not only employs a hard assignment procedure into the pooling operation but also considers the attention-based relationships between the nodes within the same cluster.
@@
Although the above hierarchical pooling methods have improved the performance of the GNN models, they still suffer from some problems when the graph is aggregated into a coarsened one. Specifically, we summarize the challenges arising in existing methods as follows. (a) Over-smoothing [25]. The node representation of each layer relies on the information propagation between neighbor nodes over the global coarsened graph structure, thus either the original nodes or the coarsened nodes can continuously propagate their information to all other nodes through the edges after multiple layers. As a result, the representations of different nodes tend to be similar or indistinguishable to each other. This phenomenon is the so-called notorious over-smoothing problem [26], seriously influencing the performance of the GNN models. (b) Degradation. This issue arises in the context of hierarchical pooling within GNNs, that adheres to deep learning architectures. When the GNN model associated with the pooling layers increases in depth, the degradation becomes a significant challenge, influencing the performance of the GNN model .
@@
The aim of this work is to overcome the aforementioned shortcomings by proposing a new pooling method, that focuses more on extracting local features from specific separated subgraphs. One key innovation of the new pooling method is that we hierarchically decompose the original graph structure into separated subgraphs, and extract the structural information from each individual subgraph. This scheme can in turn isolate the structural information between different subgraph structures, naturally providing an elegant way to reduce the over-smooth problem. Overall, the main contributions of this work are threefold.
To this end, we hierarchically decompose the global graph structure into a family of separated local subgraphs, and employ several individual local convolution units to compress each subgraph into a coarsened node. Since each local convolution units are restricted in each subgraph and the node information cannot be propagated between different subgraphs, the new pooling can significantly reduce the over-smoothing problem and provide discriminative coarsened node representations for the resulting coarsened graphs.
@@
First, we propose a novel local graph pooling method, namely the Separated Subgraph-based Hierarchical Pooling (SSHPool), to learn effective graph representations. We commence by assigning the nodes of each graph into different clusters, and this process results in a family of separated subgraphs where each subgraph consists of the nodes belonging to the same cluster and remains the edge connection between these nodes from the original graph. With these separated graphs to hand, we individually employ several local graph convolution units to compress each subgraph into a coarsened node, further transforming the original graph into a coarsened graph. Since each local convolution unit is restricted in each subgraph and the node information cannot be propagated between different subgraphs, the new pooling method can significantly reduce the over-smoothing problem and provide discriminative coarsened node representations for the resulting coarsened graphs. Furthermore, by hierarchically performing the proposed procedures on the resulting coarsened graph, the proposed SSHPool can effectively extract the hierarchical global feature of the original graph structure, encapsulating rich intrinsic structural characteristics.
@@
Second, we develop an end-to-end GNN-based framework associated with the proposed SSHPool module for graph classification. Since the proposed SSHPool consists of hierarchical substructure sets, there are multiple layer-wise stacked local convolution unit sets associated with the hierarchical substructures, i.e., the SSHPool module may be deeper. To overcome the possible degradation problem that maybe caused by the deeper hierarchical structures, we propose an attention layer for the proposed GNN-based model to accomplishes the interaction between the global features extracted from the SSHPool and the initial embeddings from the original graph, further improving the effectiveness of the proposed GNN-based model.
@@
Third, we empirically evaluate the classification performance of the GNN-based model associated with the proposed SSHPool module. Experimental results on five standard graph datasets demonstrate the effectiveness of the proposed SSHPool our model.
@@
The reminder of our paper is organized as follows. In Section II, we first introduce related work about our research. Section III presents our proposed approach and model in detail. Section IV expresses the performance of our model on real-world datasets, and also analyzes the reason for the effectiveness of it. Finally, we conclude this paper in Section V.
@@ II-A The Graph Neural Networks
GNNs have been widely employed for various graph-based applications [27, 28, 29]. In these instances, the input data is represented as a graph structure, and the aim of the GNN is to learn effective representations that can better describe the structural characteristics of graphs.
@@
where $Z_{G}$ is the extracted graph embedding vector matrix (i.e., the graph feature representation), and $\Theta$ is the set of training parameters for the GNN model. With these graph representations to hand, the downstream classifier can be employed to predict their corresponding labels.
Inspired by the traditional CNNs, the GNN methods have rapidly evolved and further form the so-called Graph Convolution Networks (GCNs) for graph structures. For instance, the GCN model proposed by Kipf et al., [6] implements the GNN model by combining the following layer-wise propagation scheme, i.e.,
@@
where $H^{(l)}\in\mathbb{R}^{n\times d}$ is the hidden embedding matrix from the $l$ layer, $W^{l}\in\mathbb{R}^{d\times d}$ is the trainable matrix for the $l$ layer, $\tilde{A}=A+I$ is the adjacency matrix associated with the self loop, $\tilde{D}=\sum_{j}\tilde{A}_{ij}$ is the corresponding degree matrix associated with $\tilde{A}$, and $H^{(l+1)}$ is the embedding matrix extracted for the next layer $l+1$ of the GNN model. Note that, $H^{(0)}$ corresponds to the node feature matrix $X$, i.e., the $H^{(0)}=X$ is the original graph feature representation. Eq.(2) indicates that each module layer of the GNN model adopts the structural message propagation function to accomplish the representation learning. Unlike traditional GNNs, the GCN is able to extract and aggregate the graph structure information within the neighborhood rooted at each node. With the increasing layers, the GCN can gradually encapsulate the high-order neighborhood information into consideration.
@@ II-B The Pooling Approaches
As one kind of information filtering and downsize method, the pooling operations are widely used in traditional CNNs. Compared to the large number of pooling strategies for traditional CNNs, the amount of graph pooling methods is fewer. Generally speaking, there are two specific types of pooling methods for graphs, namely the global pooling and the local pooling. Specifically, the global pooling approaches tend to employ some simple operations to directly compute the global representations of graphs, such as summing up or averaging all the node representations over the whole graph [24]. Since the simple summing or averaging operation cannot identifies the feature distribution of different nodes, the global pooling approaches usually have poor performance.
On the other hand, the local pooling approaches tend to gradually form a family of coarsening graph with shrinking sizes to extract global features of graph structures, they can thus reflect rich local intrinsic structural information. As a typical instance of local pooling methods, the DiffPool [17] operation is proposed to learn hierarchical graph representations. Specifically, given the input graph adjacent matrix $A^{(l)}$ and the node embedding matrix $X^{(l)}\in\mathbb{R}^{n_{l}\times d}$ for layer $l$, the DiffPool layer coarsens the input graph as follows, i.e.,
@@
where $A^{(l+1)}\in\mathbb{R}^{n_{l+1}\times d}$ and $X^{(l+1)}$ are the corresponding adjacency matrix and the node embedding matrix of the resulting coarsened graph for the next layer $l+1$, and $\mathrm{DiffPool}_{l}(\cdot)$ is the pooling function in layer $l$. Since the DiffPool method can hierarchically learn node assignment matrices that can assign the nodes to different clusters, this pooling method can aggregate the graph of the current layer into a coarsened graph for the next layer by compressing the original nodes (i.e., $X^{(l)}$) belonging to the same cluster into new coarsened nodes (i.e., $X^{(l+1)}$). Thus, the node number $n_{l+1}$ of $X^{(l+1)}$ is lower than that $n_{l}$ of $A^{(l+1)}$.
@@
Based on the basic hierarchical scheme of the DiffPool method, a number of extended hierarchical pooling methods have been developed by redesigning the assignment procedure or the coarsening procedure [30, 18, 21, 31, 32, 33]. Specifically, these hierarchical pooling methods can be divided into the following categories, by considering whether the pooling operation needs to reserve all nodes.
@@
(a) Hierarchical pooling with partial nodes. One typical idea is the adaptive structure aware pooling (ASAPool) [18] proposed by Ranjan et al. After the assignment process, the ASAPool scores the different clusters in the graph. The top scored clusters will be selected and compressed into coarsened nodes. The other clusters considered as the redundant information are abandoned. As a result, the ASAPool operation relies on partial nodes. (b) Hierarchical pooling with all nodes. One typical example is the attention-based differentiable pooling (ABDPool) [17]. Unlike the ASAPool, it tend to obtain the structural information from all node clusters.
@@
Remarks: In a summary, the pooling operation with partial nodes focuses on capturing the structural information from the key clusters, while the pooling operation with all nodes tends to encapsulate the structural information over the whole graph structure. Although these pooling significantly improve the existing GNN models, most existing hierarchical pooling still suffer from two drawbacks, that is the over-smoothing problem as well as the degradation problem. The first problem is due to the fact that the node representation of each hierarchical layer relies on the information propagation between neighbor nodes, and each node information will be propagated to all other nodes through the paths between the nodes after multiple layers, resulting in similar or indistinctive node representations for all nodes. The second problem is due to the fact that these hierarchical pooling operations are usually associated with GNNs that usually have deep computational architectures, significantly resulting in the degradation.
III Methodology

### v2

@@ Abstract
In this paper, we develop a novel local graph pooling method, namely the Separated Subgraph-based Hierarchical Pooling (SSHPool), for graph classification. We commence by assigning the nodes of a sample graph into different clusters, resulting in a family of separated subgraphs. We individually employ the local graph convolution units as the local structure to further compress each subgraph into a coarsened node, transforming the original graph into a coarsened graph. Since these subgraphs are separated by different clusters and the structural information cannot be propagated between them, the local convolution operation can significantly avoid the over-smoothing problem caused by message passing through edges in most existing Graph Neural Networks (GNNs). By hierarchically performing the proposed procedures on the resulting coarsened graph, the proposed SSHPool can effectively extract the hierarchical global features of the original graph structure, encapsulating rich intrinsic structural characteristics. Furthermore, we develop an end-to-end GNN framework associated with the SSHPool module for graph classification. Experimental results demonstrate the superior performance of the proposed model on real-world datasets.
@@ I Introduction
In recent years, deep learning methods such as the Convolutional Neural Network (CNN) have experienced a rapid developing process. Many effective CNN-based models have been applied to multiple research fields, including image segmentation [1], image classification [2], natural language processing [3], etc. The great success of the CNN encourages researchers to further explore employing it on graph structure data [4], that usually has a complex structure and is a typical instance of irregular data in non-Euclidean space. Unlike the regular data such as images and vectors, the graph data requires more effective models, that can not only analyze the attributes of the elements in the graph structure but also figure out the structural relationships between the elements. Thus, it is difficult to directly employ the traditional CNNs for analyzing the graph data.
I-A Literature Review
@@
To effectively deal with the graph data, some researchers have tried to enhance the ability of CNN models for irregular grid structures of graphs. Meanwhile, the Graph Neural Networks (GNNs) [5] are proposed to accomplish graph-based tasks. By integrating the CNN and the GNN methods together, the Graph Convolutional Network (GCN) [6] has been rapidly developed, and its related approaches [7, 8] have achieved outstanding performance in capturing the effective features of the graph data. For example, the Deep Graph Convolution Neural Network (DGCNN) [8] has leveraged the graph convolution layers with the SortPooling module to learn semantic characteristics. Based on the DGCNN models, some researchers further explore its application for different downstream tasks [9, 10, 11]. To accomplish the classification task, the Hierarchical Graph Convolutional Network (HGCN) [12] has adopted the hierarchical graph convolution to aggregate structurally similar nodes, exhibiting its strong empirical performance on some specific downstream tasks (e.g., node classification). Furthermore, by associating with the attention mechanism, some derivative GNN models, such as the Graph Transformer Network (GTN) [13] and the Spatio-Temporal Graph Transformer Network (STGTN) [14], have been developed and further improved the capability of extracting complicated graph features. These extended GNN-based models provide powerful tools for graph data analysis. However, when facing the graph representation learning task, there is usually additional requirements for these aforementioned models, i.e., they need to abstract features from original graphs with less redundant information [15]. To this end, the downsampled filtering method, namely the graph pooling operation, is highlighted.
@@
Similarly to the traditional pooling operation defined for the CNN models, the graph pooling operation proposed for the GNN models plays an important role in the sampling process. The main purpose of the graph pooling is to downsize the graph structure and make the representation learning efficient [16]. For existing works, the hierarchical scheme in the GNN models has significantly affected the definitions of novel graph pooling operations, and the researchers have focused more on the enhancement of the GNN models associated with the idea of hierarchical pooling [17, 18, 19, 20, 21]. Specifically, for the multi-layer framework of the GNN models, the original input graph can be gradually compressed into a family of coarsened graphs with shrinking sizes, so that the global feature of the original graph can be obtained through the coarsened graphs. According to the arrangement of nodes involved in the node assignment, there are two specific types of pooling methods for graphs, namely the global pooling and the local pooling. The descriptions of both the methods are shown in Figure 1.
@@
Figure 1: Descriptions of two pooling methods. Both the global and the local pooling methods convert the original input graph into the global feature. On the left side, the global pooling tends to directly sum up or average the representations of all nodes over the whole graph structure. By contrast, on the right side, the local pooling tends to assigns the nodes into different clusters, and then hierarchically aggregates the nodes from the same cluster into a coarsened node, resulting in a family of layer-wise coarsened graphs.
@@
Generally speaking, the global pooling is a common methodology in the pooling operation [22, 23]. Global operations, such as the summing pooling and the averaging pooling, can directly compute the global representation of a graph by simply summing up or averaging all the node representations over the whole graph [24]. However, since the global pooling neglects the feature distribution associated with different nodes, it cannot identify the structural differences between the nodes. Thus, the global pooling can not effectively learn the topological information residing on the local nodes and may be unsuitable for a hierarchical structure. To overcome the shortcoming, the local pooling that focuses more on the features of local nodes is developed. With this strategy, a number of local pooling operations, namely the hierarchical pooling operations, based on hierarchical structures are developed. One typical approach of them is the DiffPool [17]. The core idea is to learn the layer-wise node assignment matrices, that assign the nodes to different clusters and then hierarchically aggregate the graph of the current layer into a coarsened graph for the next layer, by compressing the nodes belonging to the same cluster into a new coarsened node. However, the associated node assignment matrix is computed based on the soft assignment procedure, i.e., each node may be simultaneously assigned to multiple clusters associated with different probabilities. For the DiffPool, the receptive field of each local patch (i.e., the cluster) may influence that of other patches. Moreover, the coarsened node representation of the DiffPool is computed by directly summing up the node representations belonging to the same cluster, the DiffPool fails to identify the differences between different nodes.
@@
Although the above hierarchical pooling methods have improved the performance of the GNN models, they still suffer from some problems when the graph is aggregated into a coarsened one. Specifically, we summarize the challenges arising in the existing methods as follows. (a) Over-smoothing [25]. The node representation of each layer relies on the information propagation between neighbor nodes over the global coarsened graph structure. Thus, either the original nodes or the coarsened nodes can continuously propagate their information to all other nodes through the edges after multiple layers. As a result, the representations of different nodes tend to be similar or indistinguishable to each other. This phenomenon is the so-called notorious over-smoothing problem [26], seriously influencing the performance of the GNN models. (b) Degradation. This issue arises in the context of the hierarchical pooling within GNNs, that adheres to deep learning architectures. When the GNN model associated with the pooling layers increases in depth, the degradation becomes a significant challenge, influencing the performance of the GNN model .
I-B Contributions
@@
The aim of our work is to overcome the aforementioned shortcomings by proposing a new pooling method, that focuses more on extracting local features from the specific separated subgraphs. One key innovation of the new pooling method is that we hierarchically decompose the original graph structure into the separated subgraphs, and extract the structural information from each individual subgraph. This scheme can in turn isolate the structural information between different subgraph structures, naturally providing an elegant way to reduce the over-smooth problem. Overall, the main contributions of this work are threefold.
@@
First, we propose a novel local graph pooling method, namely the Separated Subgraph-based Hierarchical Pooling (SSHPool), to learn effective graph representations. We commence by assigning the nodes of each graph into different clusters, and this process results in a family of separated subgraphs where each subgraph consists of the nodes belonging to the same cluster and retains the edge connection between these nodes from the original graph. With these separated graphs to hand, we individually employ several local graph convolution units to compress each subgraph into a coarsened node, further transforming the original graph into a coarsened graph. Since each local convolution unit is restricted in each subgraph and the node information cannot be propagated between different subgraphs, our new pooling method can significantly reduce the over-smoothing problem and provide discriminative coarsened node representations for the resulting coarsened graphs. Furthermore, by hierarchically performing the proposed procedures on the resulting coarsened graph, the proposed SSHPool can effectively extract the hierarchical global feature of the original graph structure, encapsulating rich intrinsic structural characteristics.
@@
Second, we develop an end-to-end GNN-based framework associated with the proposed SSHPool module for graph classification. Since the proposed SSHPool consists of hierarchical substructure sets, there are multiple layer-wise stacked local convolution unit sets associated with the hierarchical substructures, i.e., the SSHPool module may be deeper. To overcome the possible degradation problem that may be caused by the deeper hierarchical structures, we propose a graph attention layer for the proposed GNN-based model to accomplish the interaction between the global features extracted from the SSHPool and the initial embeddings from the original graph, further improving the effectiveness of the proposed GNN-based model.
@@
Third, we empirically evaluate the classification performance of our GNN model associated with the proposed SSHPool module. Experimental results on standard graph datasets demonstrate the effectiveness.
@@
This paper is organized as follows. Section II introduces the related work about our research. Section III presents the proposed approach in details. Section IV expresses the performance of our model on real-world datasets, and analyzes the reason for the effectiveness. Section V concludes this paper.
@@ II-A The Graph Neural Networks
GNNs have been widely employed for various graph-based applications [27, 28, 29]. In these instances, the input data is represented as a graph structure, and the GNN aims to learn effective representations that can better describe the structural characteristics of graphs.
@@
where $Z_{G}$ is the extracted graph embedding matrix (i.e., the graph feature representation), and $\Theta$ is the set of training parameters for the GNN model. With these graph representations to hand, the downstream classifier can be employed to predict their corresponding labels.
Inspired by the traditional CNNs, the GNN methods have rapidly evolved and further formed the so-called Graph Convolution Networks (GCNs) for graph structures. For instance, the GCN model proposed by Kipf et al., [6] implements the GNN model by combining the following layer-wise propagation scheme, i.e.,
@@
where $H^{(l)}\in\mathbb{R}^{n\times d}$ is the hidden embedding matrix from the $l$ layer, $W^{l}\in\mathbb{R}^{d\times d}$ is the trainable matrix for the $l$ layer, $\tilde{A}=A+I$ is the adjacency matrix associated with the self loop, $\tilde{D}=\sum_{j}\tilde{A}_{ij}$ is the corresponding degree matrix associated with $\tilde{A}$, and $H^{(l+1)}$ is the embedding matrix extracted for the next layer $l+1$ of the GNN model. Note that, $H^{(0)}$ corresponds to the node feature matrix $X$, i.e., the $H^{(0)}=X$ is the original graph feature representation. Eq.(2) indicates that each module layer of the GNN model adopts the structural message propagation function to accomplish the representation learning. Unlike the traditional GNNs, the GCN is able to extract and aggregate the graph structure information within the neighborhood rooted at a node. With the increasing layers, the GCN can gradually encapsulate the high-order neighborhood information into the consideration.
@@ II-B The Pooling Approaches
As a kind of information filtering and downsizing method, the pooling operations are widely used in the traditional CNNs. Compared to the large number of pooling strategies for the traditional CNNs, the amount of graph pooling methods is fewer. Generally speaking, there are two specific types of pooling methods for graphs, namely the global pooling and the local pooling. Specifically, the global pooling approaches tend to employ some simple operations to directly compute the global representations of graphs, such as summing up or averaging all the node representations over the whole graph [24]. Since the simple summing or averaging operation cannot identify the feature distribution of different nodes, the global pooling approaches usually have poor performance.
On the other hand, the local pooling approaches tend to gradually form a family coarsening graph with shrinking sizes to extract global features of graph structures, they can thus reflect rich local intrinsic structural information. As a typical instance of local pooling methods, the DiffPool [17] operation is proposed to learn hierarchical graph representations. Specifically, given the input graph adjacent matrix $A^{(l)}\in\mathbb{R}^{n_{l}\times n_{l}}$ and the node embedding matrix $X^{(l)}\in\mathbb{R}^{n_{l}\times d}$ for the layer $l$, the DiffPool layer coarsens the input graph as
@@
where $A^{(l+1)}$ and $X^{(l+1)}$ are the corresponding adjacency matrix and the node embedding matrix of the resulting coarsened graph for the next layer $l+1$, and $\mathrm{DiffPool}_{l}(\cdot)$ is the pooling function in the layer $l$. Since the DiffPool method can hierarchically learn node assignment matrices that assign the nodes to different clusters, this pooling method can aggregate the graph of the current layer into a coarsened graph for the next layer by compressing the original nodes (i.e., $X^{(l)}$) belong to the same cluster into new coarsened nodes (i.e., $X^{(l+1)}$). Thus, the node number $n_{l+1}$ of $X^{(l+1)}$ is lower than that $n_{l}$ of $X^{(l)}$.
@@
Based on the basic hierarchical scheme of the DiffPool method, a number of extended hierarchical pooling methods have been developed by redesigning the assignment procedure or the coarsening procedure [30, 18, 21, 31, 32, 33]. Specifically, these hierarchical pooling methods could be divided into the following categories, by considering whether the pooling operation needs to reserve all nodes.
@@
(a) Hierarchical pooling with partial nodes. One typical idea is the Adaptive Structure Aware Pooling (ASAPool) [18]. After the assignment process, the ASAPool scores the different clusters in the graph. The top scored clusters will be selected and compressed into coarsened nodes. The other clusters which are considered as the redundant information are abandoned. As a result, the ASAPool operation relies on the partial nodes. Similar approaches also include the Subgraph Neural Network with the Reinforcement Pooling (SUGAR) [19] that samples the subgraphs from the original graph, and adopts the top $k$ strategies to select the necessary subgraphs, dropping partial nodes during the pooling process. (b) Hierarchical pooling with all nodes. One typical example is the attention-based differentiable pooling (ABDPool) [17]. Unlike the ASAPool, it tends to obtain the structural information from all node clusters.
@@
Remarks: In summary, the pooling operation with partial nodes focuses on capturing the structural information from the key targets (e.g. node clusters, subgraphs), while the pooling operation with all nodes tends to encapsulate the structural information over the whole graph structure. Although these pooling methods significantly improve the existing GNN models, most existing hierarchical pooling still suffer from two drawbacks, that is the over-smoothing problem as well as the degradation problem. The first problem is due to the fact that the node representation of each hierarchical layer relies on the information propagation between neighbor nodes. The information of each node will be propagated to all other nodes through the paths between the nodes after multiple layers, resulting in similar or indistinctive node representations for all nodes. The second problem is due to the fact that these hierarchical pooling operations are usually associated with GNNs that typically have deep computational architectures, significantly resulting in degradation. Both problems limit the performance of the GNN-based models.
III The Methodology